OpenStudy (anonymous):

x/3+2x/15< less than with aline under it 2/3 what is the solution set

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

Less than what?

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

\[\frac{x}{3}+\frac{2x}{15}<\frac{2}{3}\] Multiply both sides by 15 first

OpenStudy (anonymous):

This will also cancel the 3s in the denominator as 15 is a multiple of 3

OpenStudy (anonymous):

yes but the symbol for less than has a line under it

OpenStudy (anonymous):

That means 'less than or equal to'. Treat it the same as a normal 'less than' algebraically.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[< \rightarrow \le\]

OpenStudy (anonymous):

ok got ya so how would i get the solution set

OpenStudy (anonymous):

Although obviously when you're computing possible answer take account that it could also be equal to the thing.

OpenStudy (anonymous):

Multiplying by 15:\[5x+2x \le 10\]

OpenStudy (anonymous):

Rearrange to make x the subject, and there is your answer.

OpenStudy (anonymous):

can you help me with that

OpenStudy (anonymous):

What is 5x+2x?

OpenStudy (anonymous):

7x

OpenStudy (anonymous):

\[7x \le 10\]

OpenStudy (anonymous):

Divide by 7 to get 1x on the left.

OpenStudy (anonymous):

\[\frac{7x}{7}=x \le\frac{10}{7}\]

OpenStudy (anonymous):

So x is equal to any number smaller than or up to 10/7

OpenStudy (anonymous):

so is the solution set x<10/7

OpenStudy (anonymous):

 Not the < sign , but with the 'or equal to' line under it, yes.